# Magnon-induced transparency of a disordered antiferromagnetic Josephson junction

### v1

@@ Abstract
The suppression of the proximity effect occurs when an AFM makes a contact with a spin-singlet superconductor. On the other hand, as noted in Brataas , the long-range proximity effect takes place for triplet Cooper pairs whose spins are perpendicular to the Néel order. In Josephson junctions (JJ) such triplets can be created by ferromagnetic interlayers placed between superconductors and AFM Malsh . A new opportunity for creating and detecting of triplet superconducting correlations in AFM opens, if one takes into account that triplets carry the angular moment which, in turn, can interact with magnetic excitations of the Néel order. In which connection, we considered the effect of magnons on the critical current through a phase-biased S-AFM-S planar junction. The idea is that magnons can transfer their angular moment to electrons. By flipping electron spins magnons promote a transformation of singlet pairing correlations into triplet ones. Due to this conversion magnetic excitations lead to a strong increase of the critical current through a disordered JJ whose length is larger than the electron’s mean free path, because triplet correlations can extend over large distances. Such a strong effect of magnons on the Josephson current opens the way for new spintronic applications in hybrid systems which combine superconducting circuits and magnetic materials.
@@
The main physics of the considered effect is explained in Fig.1. The dc current is driven by the phase bias. Two-magnon absorption-emission processes are taken into account perturbatively. For simplicity, the coherent excitation of the Néel order is treated as a classical magnetic wave whose length is much larger than the distance between S-contacts. Magnons may be excited and transmitted to JJ by many ways. Here, it is assumed that the magnetic excitation is produced by the spin-orbit torque oscillator Cheng ; Khymyn ; Khymyn2 . The corresponding setup is discussed below in Supplemental material (SM).
@@
The observable current can be calculated by using the disorder averaged magnon-excited nonequilibrium Green function $\mathcal{G}^{\lambda}_{\mathbf{k}R(L)}(\omega,\mathbf{r})$, which takes account of the interaction with superconductor’s functions $G^{s}_{\mathbf{k}^{\prime}L(R)}$. This interaction is calculated within the second order perturbation theory with respect to $t^{\lambda}_{R(L)\mathbf{k},\mathbf{k}^{\prime}}$. The subscripts $R$ and $L$ in $\mathcal{G}$ indicate that this function is calculated by taking into account the electron tunneling from the right, or left superconductors, respectively. Green’s functions in AFM and S are represented by $\times 2$ Keldysh matrices Rammer given by $G_{11}=G^{\mathrm{r}},G_{22}=G^{\mathrm{a}},G_{12}=G^{\mathrm{K}}$ and $G_{21}=0$, where the superscripts $\mathrm{r},\mathrm{a}$ and $\mathrm{K}$ denote retarded, advanced and Keldysh functions, respectively. So, by expanding the average of Eq.(2) with respect to $t^{\lambda}_{R(L)\mathbf{k},\mathbf{k}^{\prime}}$ the stationary electric current can be written in the form
@@
where the trace is taken over spin and Nambu variables, while the superscript $\mathrm{K}$ denotes the Keldysh component of the Green’s functions product. In the leading approximation with respect to the tunneling amplitude $G^{s\mathrm{K}}$ is given by the equilibrium function $G^{s\mathrm{K}}=(G^{s\mathrm{r}}-G^{s\mathrm{a}})\tanh(\omega/2T)$, while $\mathcal{G}^{\lambda\mathrm{K}}_{\mathbf{k}\mathbf{q}R(L)}$ is perturbed by magnons from the thermodynamic equilibrium. In the spatial representation the functions $w^{L(R)}$ describe the shapes of S-AFM interface contacts. They turn to zero outside the interfaces which have the rectangular form. For identical contacts of the width $w$, which are placed at the distance $l$ from each other, the Fourier components of these functions are given by $w^{L}_{\mathbf{q}}=w^{R}_{\mathbf{q}}\equiv w_{\mathbf{q}}=\exp(iq_{x}l)q^{-1}_{x}\sin wq_{x}$. The wavevector $q$ in Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) is assumed to be much less than vectors $k$ and $k^{\prime}$ which are close to respective Fermi surfaces of S and AFM. In the case of a disordered interface it is reasonable to simplify Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) by ignoring the dependence of $|t^{\lambda}_{L(R)\mathbf{k},\mathbf{k}^{\prime}}|^{2}$ on $\mathbf{k}$ and $\mathbf{k}^{\prime}$. Moreover, the tunneling amplitude is assumed to be independent on the sublattice index $\lambda$. Therefore, below we set $|t^{\lambda}_{L\mathbf{k},\mathbf{k}^{\prime}}|^{2}=|t^{\lambda}_{R\mathbf{k},\mathbf{k}^{\prime}}|^{2}=|t|^{2}$. As a consequence, the interaction with superconducting contacts leads to the selfenergy of $\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}R(L)}(\omega,\mathbf{r})$ in the form
@@
The unperturbed impurity averaged Green function $G_{\mathbf{k}}(\omega)$ of the AFM is given by the tight binding model on a bipartite cubic lattice. Néel-ordered localized spins $S$ and conduction electrons interact through the exchange interaction $SJ\sigma_{z}\lambda_{z}$. This model results in the electron band energies ±$\pm E_{\mathbf{k}}$ where $E_{\mathbf{k}}=(\epsilon^{2}_{\mathbf{k}}+S^{2}J^{2})^{1/2}$ and $\epsilon_{\mathbf{k}}=-t(\cos ak_{x}+\cos ak_{y}+\cos ak_{z})$, with $t$ denoting the electron’s hopping constant between nearest sites of the lattice. The disorder is created by random fluctuations $u_{i}$ of electron energies on lattice cites. These fluctuations are not correlated. Therefore, $\overline{u_{i}u_{j}}=\overline{u^{2}}\delta_{ij}$. The disorder gives rise to the finite mean elastic scattering time $1/2\tilde{\Gamma}$ of electrons, where $\tilde{\Gamma}=(1+\xi^{2})\Gamma,\Gamma=\pi\overline{u^{2}}N_{\mu}/2$, $\mu$ is the chemical potential which is counted from the middle of the band, $\xi=SJ/\mu$, and $N_{\mu}$ is the state density at the Fermi level, per a unit cell. So, the retarded and advanced components of $G_{\mathbf{k}}(\omega)$ can be written, as Malshukov2
@@
Besides the interaction with magnons, the expression for $\mathcal{G}(\omega)$ includes selfenergy Eq.(4) which is associated with the S-AFM coupling at contacts. The interaction with contacts results in the Andreev reflection of electrons from them. In turn, the impurity averaging gives rise to the coherent scattering of electrons and Andreev-reflected holes on the same impurities. This process can be described by diffusion motion of the so correlated pairs between contacts Aslamazov , as shown in Fig.1. Such a diffusion takes place at $\Gamma\gg\omega,\Omega,Dq^{2}$, where $D=\overline{v_{F}^{2}}/4\Gamma$ is the diffusion constant and $\overline{v_{F}^{2}}$ denotes averaging over the Fermi surface. During random motion electrons can flip their spins by interacting with magnons, as seen from Eqs.(7-8). Therefore, the total spin of correlated pairs may be different from zero, which is originally provided by the proximity to the singlet superconductor. In contrast to singlets, triplet Cooper correlations propagate over much larger distances. Formally, such diffusion processes can be expressed in terms of the diffusion propagators $\mathcal{D}_{00}^{\mathrm{r}(\mathrm{a})}(\omega)$ and $\mathcal{D}_{ij}^{\mathrm{r}(\mathrm{a})}(\omega)$ where $(i,j)\in(x,y)$ In the absence of the spin-orbit interaction these functions are given by Malsh
@@
where $d^{11(22)}_{xx(yy)}=-\Gamma(1+\xi^{2})$, $d^{12(21)}_{xx(yy)}=-\Gamma(1-\xi^{2})$, $d^{11(22)}_{xy}=-d^{11(22)}_{yx}=\pm 2i\Gamma\xi$, $d^{12(21)}_{xy}=0$, and $d_{0}=-\Gamma(1-\xi^{2})$. The diffusion coefficients are renormalized, as $D_{0}=(1-\xi^{2})(1+\xi^{2})^{-2}D$ and $D_{\perp}=(1+\xi^{2})^{-1}D$. Expressions for advanced functions $\mathcal{D}^{\mathrm{a}}(\omega,\mathbf{q})$ are represented by Eq.(9) with complex conjugated denominators. One can see that the spatial behavior of $\mathcal{D}^{\mathrm{r}}_{ij}(\omega,q_{x})$ is determined by its pole at $q_{x}\simeq(2i\omega/D)^{1/2}$. Since $\omega\sim\Delta_{R(L)}$, this pole corresponds to the long-range diffusion of triplet correlations, on distances comparable to a paramagnetic metal. In contrast, the diffusion of singlets, which is given by $\mathcal{D}^{\mathrm{r}}_{00}$, decreases much faster when $\xi\sim 1$, at the length $|q_{x}|^{-1}\sim(D_{0}/\Gamma)^{1/2}$. It is important that long-range functions $\mathcal{D}_{ij}(\omega,\mathbf{q})$ appear only in two-magnon processes. They are absent in the linear perturbation theory which might lead to the time-dependent current. In the latter case the diffusion occurs through propagators $\mathcal{D}_{0i}$ and $\mathcal{D}_{i0}$. In the considered model of AFM they are zero, even if the spin-orbit interaction is taken into account Malsh .
Green function-The total AFM Green functions $\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}}(\omega)$ can be obtained by combining diffusion propagators from Eq.(9) with selfenergy Eq.(4) and Green’s functions given by Eqs.(6-8). It should be taken into account that, if $|t^{\lambda}_{L(R)\mathbf{k},\mathbf{k}^{\prime}}|^{2}$ does not depend on wavevectors and sublattice variables, the AFM Green’s function enters in Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) as the sum $\sum_{\mathbf{k}\lambda}\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}}$. Most simple expressions may be written for $\mathcal{G}^{\mathrm{r}(\mathrm{a})}_{\mathbf{k}\mathbf{q}}$, because the latter involves only retarded (advanced) functions from Eq.(7), while $\mathcal{G}^{K}$ contains products of $G^{\mathrm{r}}$ and $G^{\mathrm{a}}$ (see SM). The retarded function enters Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) as
@@
Numerical results- Fig.2 depicts the dependence of the normalized critical current on the frequency of evanescent magnetic waves which are excited in AFM due to its contact with an antiferromagnetic nanooscillator (see the corresponding set up in Fig.S2 in SM). We set $\Omega_{0}=4\Delta$. Therefore, at $\Delta=$1meV $\Omega_{0}$ is about 1THz. Only the magnon-induced part of the current is shown at Fig.2, while the unperturbed singlet current is very small in the range of $l>l_{c}$. For instance, at $l=l_{c}$, $\xi=0.5$, and $k_{B}T=0.01\Delta$ the normalized singlet current is only 6$\cdot$10-3, that is much less than the current in Fig.2 at $\Omega=1.8\Delta$. As seen from Fig.2, the magnon effect does not considerably vary with the temperature. At the same time, its frequency dependence is rather nontrivial. The critical current changes sign near $\Omega=1.5\Delta$. Such a 0-$\pi$ transition is also takes place at the varying junction length, as seen in Fig.3. It is important that the frequency of spin precession generated by a nanooscillator in Ref.Khymyn can be widely varied by the electric current in the Pt film. Hence, in this way one can also control the Josephson current and switch the JJ into the 0-$\pi$ transition regime.
@@
$j_{L(R)}=e|t|^{2}\sum_{\mathbf{k},\mathbf{q}\lambda}w^{L}_{\mathbf{q}}w^{R}_{\mathbf{q}}\times$
$\int\frac{d\omega}{2\pi}\mathrm{Tr}\left([\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}R(L)}(\omega),\Sigma_{L(R)}(\omega)]\tau_{3}\right)^{\mathrm{K}}\,,$ (S1)
where selfenergy $\Sigma_{R(L)}(\omega)$ is given by Eqs.(4,5), with $G^{s}_{R(L)\mathbf{k}}(\omega)$ represented by the Keldysh matrix. The AFM Green function in Eq.(I.1) should be averaged over disorder. In the Born approximation, which is valid when the electron’s Fermi wave length is much smaller than the mean free path, the averaging procedure is reduced to the calculation of the so called ladder perturbation series Rammer , as shown in Fig.(S1). Elements of the ladders are represented by two-particle loops which were calculated in Ref.Malsh for a thermodynamically equilibrium AFM weak link. In the considered here nonequilibrium case these two-particle functions have the form
$\Pi^{\mathrm{r}(\mathrm{a}),\lambda\lambda^{\prime}}_{ij,12,\mathbf{q}}(\omega+\omega^{\prime})=$
$\frac{u^{2}}{2}\sum_{\mathbf{k}}\mathrm{Tr}[G^{\mathrm{r}(\mathrm{a}),\lambda\lambda^{\prime}}_{11,\mathbf{k}+\mathbf{q}}(\omega)\sigma^{j}G^{\mathrm{r}(\mathrm{a}),\lambda^{\prime}\lambda}_{22\mathbf{k}}(\omega^{\prime})\sigma^{i}]\,,$ (S2)
@@
$\Pi^{\mathrm{r}(\mathrm{a}),\lambda,\pm}_{0j,12,\mathbf{q}}(\omega,\Omega)=\frac{u^{2}}{2}\times$
$\sum_{\mathbf{k},\lambda^{\prime}}\mathrm{Tr}[\sigma^{0}G^{(1)\mathrm{r}(\mathrm{a}),\lambda^{\prime}\lambda}_{11,\mathbf{k}+\mathbf{q}}(\omega,\omega\pm\Omega)\sigma^{j}G^{\mathrm{r}(\mathrm{a}),\lambda\lambda^{\prime}}_{22\mathbf{k}}(\omega)]\,$ (S3)
@@
$\Pi^{\mathrm{raa}\lambda,\pm}_{0j,12,\mathbf{q}}(\omega,\Omega)=\frac{u^{2}}{2}\times$
$\sum_{\mathbf{k},\lambda^{\prime}}\mathrm{Tr}[\sigma^{0}G^{(1),\lambda^{\prime}\lambda}_{11,\mathbf{k}+\mathbf{q}}(\omega,\omega\pm\Omega)\sigma^{j}G^{\mathrm{a},\lambda\lambda^{\prime}}_{22\mathbf{k}}(\omega)]\,,$ (S4)
where $j\in(x,y)$, $\sigma^{0}$ is a unit matrix, and the functions $G^{(1)}$ in Eq.(I.1) and Eq.(I.1) are given by Eq.(7) and Eq.(8), respectively. There are also functions $\Pi^{\mathrm{r}(\mathrm{a}),\lambda,\pm}_{j0,12,\mathbf{q}}$ and $\Pi^{\mathrm{raa}\lambda,\pm}_{j0,12,\mathbf{q}}$, where $\sigma^{0}$ and $\sigma^{j}$ have exchanged places. Other fermion loops can be obtained by interchanging positions of $G$ and $G^{(1)}$ in Eq.(I.1) and Eq.(I.1). Such functions will be denoted as $\overline{\Pi}$. It is important that magnons always involve spin flipping of electrons. Therefore, the correlation functions given by Eq.(I.1) and Eq.(I.1), as well as other correlation functions which are associated with magnons, are responsible for a singlet-triplet mixing which leads to $0j$ and $j0$ matrix elements of $\Pi$. On this reason in Eq.(I.1) and Eq.(I.1) there is the summation on the sublattice index $\lambda^{\prime}$. That is because $\Pi_{00}$ in Eq.(I.1) does not depend on the sublattice variable. Therefore the neighbors of this singlet loop in ladder series can be independently summed up on this variable.
Eqs.(I.1-I.1) can be simplified by ignoring their dependence on $\omega,\Omega$ and $q$, because within the diffusion approximation the latter are much smaller than $\Gamma$. As a result, we arrive to a set of matrices which conveniently represent Eq.(I.1), Eq.(I.1) and other similar correlators. For instance, $\Pi^{\mathrm{r},\lambda,\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{rr}\lambda,\pm}_{\mathrm{r}0i}$ and $\overline{\Pi}^{\mathrm{r},\lambda,\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{r}\lambda\pm}_{\mathrm{rr}0i}$. The difference between these two examples is that they are given by the Green functions products $G^{(1)}_{11}G_{22}$ and $G_{11}G^{(1)}_{22}$, respectively. Other examples are $\Pi^{\mathrm{raa},\lambda\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{ra}\lambda\pm}_{\mathrm{a}0i}$ and $\overline{\Pi}^{\mathrm{rra},\lambda\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{r}\lambda\pm}_{\mathrm{ra}0i}$.
@@
$S_{\mathbf{q}0}(\omega)=[\mathcal{D}^{\mathrm{r}}_{00}(\omega,\mathbf{q})\Sigma^{\mathrm{r}}(\omega)-$
$\mathcal{D}^{\mathrm{a}}_{00}(\omega,\mathbf{q})\Sigma^{\mathrm{a}}(\omega)]\tanh\frac{\omega}{2k_{B}T}\,$ (S6)
@@
where $\overline{\nu}=-\nu$ and $\mathfrak{R}$ is the vertex renormalization constant which appears in Eq.(8). Its calculation is presented in the next subsection. The matrices $\Pi$ in Eqs.(I.1-I.1) can be calculated from Eq.(I.1),Eq.(I.1) and Eqs.(6-8) as

### v2

@@ Abstract
The suppression of the proximity effect occurs when an AFM makes a contact with a spin-singlet superconductor. On the other hand, as noted in Brataas , the long-range proximity effect takes place for triplet Cooper pairs whose spins are perpendicular to the Néel order. Such a penomenon is well known for ferromagnets with an inhomogeneous magnetization Bergeret2 ; Kadigrobov . In Josephson junctions (JJ) such triplets can be created by ferromagnetic interlayers placed between superconductors and AFM Malsh . Other ways for creating the triplet pairing in proximized FM and AFM were considered in Linder ; Rezaei ; Chaou . A new opportunity for creating and detecting of triplet superconducting correlations in AFM opens, if one takes into account that triplets carry the angular moment which, in turn, can interact with magnetic excitations of the Néel order. In which connection, we considered the effect of magnons on the critical current through a phase-biased S-AFM-S planar junction. The idea is that magnons can transfer their angular moment to electrons. By flipping electron spins magnons promote a transformation of singlet pairing correlations into triplet ones. Due to this conversion magnetic excitations lead to a strong increase of the critical current through a disordered JJ whose length is larger than the electron’s mean free path, because triplet correlations can extend over large distances. Such a strong effect of magnons on the Josephson current opens the way for new spintronic applications in hybrid systems which combine superconducting circuits and magnetic materials.
@@
The main physics of the considered effect is explained in Fig.1. The dc current is driven by the phase bias. Two-magnon absorption-emission processes are taken into account perturbatively. For simplicity, the coherent excitation of the Néel order is treated as a classical magnetic wave whose length is much larger than the distance between S-contacts. Magnons may be excited and transmitted to JJ by many ways. Here, it is assumed that the magnetic excitation is produced by the spin-orbit torque oscillator Cheng ; Khymyn ; Khymyn2 ; Liu . The corresponding setup is discussed below in Supplemental material (SM).
@@
The observable current can be calculated by using the disorder averaged magnon-excited nonequilibrium Green function $\mathcal{G}^{\lambda}_{\mathbf{k}R(L)}(\omega,\mathbf{r})$, which takes account of the interaction with superconductor’s functions $G^{s}_{\mathbf{k}^{\prime}L(R)}$. This interaction is calculated within the second order perturbation theory with respect to $t^{\lambda}_{R(L)\mathbf{k},\mathbf{k}^{\prime}}$. The subscripts $R$ and $L$ in $\mathcal{G}$ indicate that this function is calculated by taking into account the electron tunneling from the right, or left superconductors, respectively. Green’s functions in AFM and S are represented by $2\times$ 2 Keldysh matrices Rammer given by $G_{11}=G^{\mathrm{r}},G_{22}=G^{\mathrm{a}},G_{12}=G^{\mathrm{K}}$ and $G_{21}=0$, where the superscripts $\mathrm{r},\mathrm{a}$ and $\mathrm{K}$ denote retarded, advanced and Keldysh functions, respectively. So, by expanding the average of Eq.(2) with respect to $t^{\lambda}_{R(L)\mathbf{k},\mathbf{k}^{\prime}}$ the stationary electric current can be written in the form
@@
where the trace is taken over spin and Nambu variables, while the superscript $\mathrm{K}$ denotes the Keldysh component of the Green’s functions product. In the leading approximation with respect to the tunneling amplitude $G^{s\mathrm{K}}$ is given by the equilibrium function $G^{s\mathrm{K}}=(G^{s\mathrm{r}}-G^{s\mathrm{a}})\tanh(\omega/2T)$, while $\mathcal{G}^{\lambda\mathrm{K}}_{\mathbf{k}\mathbf{q}R(L)}$ is perturbed by magnons from the thermodynamic equilibrium. In the spatial representation the functions $w^{L(R)}$ describe the shapes of S-AFM interface contacts. They turn to zero outside the interfaces which have the rectangular form. For identical contacts of the width $w$, which are placed at the distance $l$ from each other, the Fourier components of these functions are given by $w^{L}_{\mathbf{q}}=w^{R}_{\mathbf{q}}\equiv w_{\mathbf{q}}=\exp(iq_{z}l)q^{-1}_{z}\sin wq_{z}$. The wavevector $q$ in Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) is assumed to be much less than vectors $k$ and $k^{\prime}$ which are close to respective Fermi surfaces of S and AFM. In the case of a disordered interface it is reasonable to simplify Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) by ignoring the dependence of $|t^{\lambda}_{L(R)\mathbf{k},\mathbf{k}^{\prime}}|^{2}$ on $\mathbf{k}$ and $\mathbf{k}^{\prime}$. Moreover, the tunneling amplitude is assumed to be independent on the sublattice index $\lambda$. Therefore, below we set $|t^{\lambda}_{L\mathbf{k},\mathbf{k}^{\prime}}|^{2}=|t^{\lambda}_{R\mathbf{k},\mathbf{k}^{\prime}}|^{2}=|t|^{2}$. As a consequence, the interaction with superconducting contacts leads to the selfenergy of $\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}R(L)}(\omega,\mathbf{r})$ in the form
@@
The unperturbed impurity averaged Green function $G_{\mathbf{k}}(\omega)$ of the AFM is given by the tight binding model on a bipartite cubic lattice. Néel-ordered localized spins $S$ and conduction electrons interact through the exchange interaction $SJ\sigma_{z}\lambda_{z}$. This model results in the electron band energies $\pm E_{\mathbf{k}}$ where $E_{\mathbf{k}}=(\epsilon^{2}_{\mathbf{k}}+S^{2}J^{2})^{1/2}$ and $\epsilon_{\mathbf{k}}=-t(\cos ak_{x}+\cos ak_{y}+\cos ak_{z})$, with $t$ denoting the electron’s hopping constant between nearest sites of the lattice. The disorder is created by random fluctuations $u_{i}$ of electron energies on lattice cites. These fluctuations are not correlated. Therefore, $\overline{u_{i}u_{j}}=\overline{u^{2}}\delta_{ij}$. The disorder gives rise to the finite mean elastic scattering time $1/2\tilde{\Gamma}$ of electrons, where $\tilde{\Gamma}=(1+\xi^{2})\Gamma,\Gamma=\pi\overline{u^{2}}N_{\mu}/2$, $\mu$ is the chemical potential which is counted from the middle of the band, $\xi=SJ/\mu$, and $N_{\mu}$ is the state density at the Fermi level, per a unit cell. So, the retarded and advanced components of $G_{\mathbf{k}}(\omega)$ can be written, as Malshukov2
@@
Besides the interaction with magnons, the expression for $\mathcal{G}(\omega)$ includes selfenergy Eq.(4) which is associated with the S-AFM coupling at contacts. The interaction with contacts results in the Andreev reflection of electrons from them. In turn, the impurity averaging gives rise to the coherent scattering of electrons and Andreev-reflected holes on the same impurities. This process can be described by diffusion motion of the so correlated pairs between contacts Aslamazov , as shown in Fig.1. Such a diffusion takes place at $\Gamma\gg\omega,\Omega,Dq^{2}$, where $D=\overline{v_{F}^{2}}/4\Gamma$ is the diffusion constant and $\overline{v_{F}^{2}}$ denotes averaging over the Fermi surface. During random motion electrons can flip their spins by interacting with magnons, as seen from Eqs.(7-8). Therefore, the total spin of correlated pairs may be different from zero, which is originally provided by the proximity to the singlet superconductor. In contrast to singlets, triplet Cooper correlations propagate over much larger distances. Formally, such diffusion processes can be expressed in terms of the diffusion propagators $\mathcal{D}_{00}^{\mathrm{r}(\mathrm{a})}(\omega)$ and $\mathcal{D}_{ij}^{\mathrm{r}(\mathrm{a})}(\omega)$ where $(i,j)\in(x,y)$ . In the absence of the spin-orbit interaction these functions are given by Malsh
@@
where $d^{11(22)}_{xx(yy)}=-\Gamma(1+\xi^{2})$, $d^{12(21)}_{xx(yy)}=-\Gamma(1-\xi^{2})$, $d^{11(22)}_{xy}=-d^{11(22)}_{yx}=\pm 2i\Gamma\xi$, $d^{12(21)}_{xy}=0$, and $d_{0}=-\Gamma(1-\xi^{2})$. The diffusion coefficients are renormalized, as $D_{0}=(1-\xi^{2})(1+\xi^{2})^{-2}D$ and $D_{\perp}=(1+\xi^{2})^{-1}D$. Expressions for advanced functions $\mathcal{D}^{\mathrm{a}}(\omega,\mathbf{q})$ are represented by Eq.(9) with complex conjugated denominators. One can see that the spatial behavior of $\mathcal{D}^{\mathrm{r}}_{ij}(\omega,q_{z})$ is determined by its pole at $q_{z}\simeq(2i\omega/D)^{1/2}$, by taking into account that $q_{x}=0$ and in a thin film, whose thickness is much less than $l$, $q_{y}=0$. Since $\omega\sim\Delta_{R(L)}$, this pole corresponds to the long-range diffusion of triplet correlations, on distances comparable to a paramagnetic metal. In contrast, the diffusion of singlets, which is given by $\mathcal{D}^{\mathrm{r}}_{00}$, decreases much faster when $\xi\sim 1$, at the length $|q_{z}|^{-1}\sim(D_{0}/\Gamma)^{1/2}$. It is important that long-range functions $\mathcal{D}_{ij}(\omega,\mathbf{q})$ appear only in two-magnon processes. They are absent in the linear perturbation theory which might lead to the time-dependent current. In the latter case the diffusion occurs through propagators $\mathcal{D}_{0i}$ and $\mathcal{D}_{i0}$. In the considered model of AFM they are zero, even if the spin-orbit interaction is taken into account Malsh .
Total Green function-The total AFM Green functions $\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}}(\omega)$ can be obtained by combining diffusion propagators from Eq.(9) with selfenergy Eq.(4) and Green’s functions given by Eqs.(6-8). It should be taken into account that, if $|t^{\lambda}_{L(R)\mathbf{k},\mathbf{k}^{\prime}}|^{2}$ does not depend on wavevectors and sublattice variables, the AFM Green’s function enters in Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) as the sum $\sum_{\mathbf{k}\lambda}\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}}$. Most simple expressions may be written for $\mathcal{G}^{\mathrm{r}(\mathrm{a})}_{\mathbf{k}\mathbf{q}}$, because the latter involves only retarded (advanced) functions from Eq.(7), while $\mathcal{G}^{K}$ contains products of $G^{\mathrm{r}}$ and $G^{\mathrm{a}}$ (see SM). The retarded function enters Eq.(Magnon-induced transparency of a disordered antiferromagnetic Josephson junction) as
@@
[…]
@@
$j_{L(R)}=e|t|^{2}\sum_{\mathbf{k},\mathbf{q}\lambda}w^{L}_{\mathbf{q}}w^{R}_{\mathbf{q}}\int\frac{d\omega}{2\pi}\mathrm{Tr}\left([\mathcal{G}^{\lambda\lambda}_{\mathbf{k}\mathbf{q}R(L)}(\omega),\Sigma_{L(R)}(\omega)]\tau_{3}\right)^{\mathrm{K}}\,$, (S1)
where selfenergy $\Sigma_{R(L)}(\omega)$ is given by Eqs.(4,5), with $G^{s}_{R(L)\mathbf{k}}(\omega)$ represented by the Keldysh matrix. The AFM Green function in Eq.(S1) should be averaged over disorder. In the Born approximation, which is valid when the electron’s Fermi wave length is much smaller than the mean free path, the averaging procedure is reduced to the calculation of the so called ladder perturbation series Rammer , as shown in Fig.(S1). Elements of the ladders are represented by two-particle loops which were calculated in Ref.Malsh for a thermodynamically equilibrium AFM weak link. In the considered here nonequilibrium case these two-particle functions have the form
$\Pi^{\mathrm{r}(\mathrm{a}),\lambda\lambda^{\prime}}_{ij,12,\mathbf{q}}(\omega+\omega^{\prime})=\frac{u^{2}}{2}\sum_{\mathbf{k}}\mathrm{Tr}[G^{\mathrm{r}(\mathrm{a}),\lambda\lambda^{\prime}}_{11,\mathbf{k}+\mathbf{q}}(\omega)\sigma^{j}G^{\mathrm{r}(\mathrm{a}),\lambda^{\prime}\lambda}_{22\mathbf{k}}(\omega^{\prime})\sigma^{i}]\,$, (S2)
@@
$\Pi^{\mathrm{r}(\mathrm{a}),\lambda,\pm}_{0j,12,\mathbf{q}}(\omega,\Omega)=\frac{u^{2}}{2}\sum_{\mathbf{k},\lambda^{\prime}}\mathrm{Tr}[\sigma^{0}G^{(1)\mathrm{r}(\mathrm{a}),\lambda^{\prime}\lambda}_{11,\mathbf{k}+\mathbf{q}}(\omega,\omega\pm\Omega)\sigma^{j}G^{\mathrm{r}(\mathrm{a}),\lambda\lambda^{\prime}}_{22\mathbf{k}}(\omega)]\,$ (S3)
@@
$\Pi^{\mathrm{raa}\lambda,\pm}_{0j,12,\mathbf{q}}(\omega,\Omega)=\frac{u^{2}}{2}\sum_{\mathbf{k},\lambda^{\prime}}\mathrm{Tr}[\sigma^{0}G^{(1),\lambda^{\prime}\lambda}_{11,\mathbf{k}+\mathbf{q}}(\omega,\omega\pm\Omega)\sigma^{j}G^{\mathrm{a},\lambda\lambda^{\prime}}_{22\mathbf{k}}(\omega)]\,$, (S4)
where $j\in(x,y)$, $\sigma^{0}$ is a unit matrix, and the functions $G^{(1)}$ in Eq.(S3) and Eq.(S4) are given by Eq.(7) and Eq.(8), respectively. There are also functions $\Pi^{\mathrm{r}(\mathrm{a}),\lambda,\pm}_{j0,12,\mathbf{q}}$ and $\Pi^{\mathrm{raa}\lambda,\pm}_{j0,12,\mathbf{q}},$ where $\sigma^{0}$ and $\sigma^{j}$ have exchanged places. Other fermion loops can be obtained by interchanging positions of $G$ and $G^{(1)}$ in Eq.(S3) and Eq.(S4). Such functions will be denoted as $\overline{\Pi}$. It is important that magnons always involve spin flipping of electrons. Therefore, the correlation functions given by Eq.(S3) and Eq.(S4), as well as other correlation functions which are associated with magnons, are responsible for a singlet-triplet mixing which leads to $0j$ and $j0$ matrix elements of $\Pi$. On this reason in Eq.(S3) and Eq.(S4) there is the summation on the sublattice index $\lambda^{\prime}$. That is because $\Pi_{00}$ in Eq.(S2) does not depend on the sublattice variable. Therefore the neighbors of this singlet loop in ladder series can be independently summed up on this variable.
Eqs.(S3-S4) can be simplified by ignoring their dependence on $\omega,\Omega$ and $q$, because within the diffusion approximation the latter are much smaller than $\Gamma$. As a result, we arrive to a set of matrices which conveniently represent Eq.(S3), Eq.(S4) and other similar correlators. For instance, $\Pi^{\mathrm{r},\lambda,\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{rr}\lambda,\pm}_{\mathrm{r}0i}$ and $\overline{\Pi}^{\mathrm{r},\lambda,\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{r}\lambda\pm}_{\mathrm{rr}0i}$. The difference between these two examples is that they are given by the Green functions products $G^{(1)}_{11}G_{22}$ and $G_{11}G^{(1)}_{22}$, respectively. Other examples are $\Pi^{\mathrm{raa},\lambda\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{ra}\lambda\pm}_{\mathrm{a}0i}$ and $\overline{\Pi}^{\mathrm{rra},\lambda\pm}_{0i,12,\mathbf{q}}(\omega,\Omega)\rightarrow\Pi^{\mathrm{r}\lambda\pm}_{\mathrm{ra}0i}$.
@@
$S_{\mathbf{q}0}(\omega)=[\mathcal{D}^{\mathrm{r}}_{00}(\omega,\mathbf{q})\Sigma^{\mathrm{r}}(\omega)-\mathcal{D}^{\mathrm{a}}_{00}(\omega,\mathbf{q})\Sigma^{\mathrm{a}}(\omega)]\tanh\frac{\omega}{2k_{B}T}\,$ (S6)
@@
where $\overline{\nu}=-\nu$ and $\mathfrak{R}$ is the vertex renormalization constant which appears in Eq.(8). Its calculation is presented in the next subsection. The matrices $\Pi$ in Eqs.(I.1-I.1) can be calculated from Eq.(S3),Eq.(S4) and Eqs.(6-8) as